If the board of directors approves a two for one stock split, an investor who owns...

  • If the board of directors approves a two for one stock split, an investor who owns 280 shares before the split owns ____________ shares after the split.

    Multiple Choice

  • 280

  • 140

  • 420

  • 560

  • 840

James Turbyfil purchased 130 shares of stock for $74. James also paid $125 commission. What was the total purchase price for this transaction?

Multiple Choice

  • $74.00

  • $9,767.50

  • $9,657.50

  • $9,745.00

  • $125.00

Jo Bower owns 170 shares of stock. She purchased the stock for $38 a share. She sold her stock for $44 a share. The commissions required to buy and sell her stock totaled $310. Assuming that she received no dividends during the time she owned the stock, what is her total return for this investment?

Multiple Choice

  • $710

  • $1,330

  • $3,800

  • $638

  • $1,020

  • Beverly Frickel purchased 180 shares of GS stock for $41.50 per share. Her commission for this purchase was $45. She sold the stock two years later for $55 per share and a commission of $60. While she held the stock it paid a dividend of $3.10 per share. What was Beverly's total dollar return on this stock?

    Multiple Choice

  • $2,988

  • $2,430

  • $2,325

  • $2,883

  • $558
Answer #1

1) 280*2 = 560

2) (130*74) = 9620+125 = 9745

3) (170*44)-(170*38)-310 = 7480-6460-310 = 1330

4) (180*55) - (180*41.5) -45-60 + (3.1*180) = 9900-7470-45-60+558 = 2883